A helium-filled weather balloon has a volume of 844 L at 14.9°C and 759 mmHg. It is released and rises to an altitude of 3.15 km, where the pressure is 586 mmHg and the temperature is –3.1°C.

The volume of the balloon at this altitude is___ L.


1
Expert's answer
2021-04-25T23:10:59-0400

Here we will apply the combined Gas Law


"\dfrac{P1V1}{T1}=\dfrac{P2V2}{T2}"


Thus;"V2=\dfrac{P1V1T2}{P2T1}"


Since

P1=756mmHg

P2=586mmHg

T1=(14.9oC + 273) = 287.9K

T2=(-3.1oC+273) = 269.9K

V1= 844L

V2?


"V2=\dfrac{756mmHg x 844L x 269.9K}{586mmHg x 287.9K} = 1024.82L"